OpenStudy (anonymous):

Find the missing value: -2/5· ? = 1 please help..

Directrix (directrix):

-2/5· ? = 1 Multiply by sides of the equation by -5/2 (-5/2 * -2/5) · ? = 1 * -5/2 1 * ? = -5/2 ? = -5/2

Directrix (directrix):

(-5/2 * -2/5) are reciprocals of each other and multiply to 1.

Directrix (directrix):

@chandlerm